OpenStudy (notamathgenius):

What would be the following system of equations looks like and augmented Matrix? x+y=5 3x+y=-1

OpenStudy (amistre64):

what does your work look like?

OpenStudy (notamathgenius):

I haven't done any, I was coming to look for help before I did anything, I don't wanna screw it up.

OpenStudy (amistre64):

you have to "screw it up" in order to learn it better

OpenStudy (amistre64):

pull off the coeffs, and run your row operations ...

OpenStudy (notamathgenius):

Awww, work darn. I don't have any. I know the answer to the X and Y is (1,4)

OpenStudy (amistre64):

the the augment will have to end in a column of 1 4

OpenStudy (notamathgenius):

None have "1,4" its all 3 -1 0 5

OpenStudy (amistre64):

assuming youve got the correct 1,4 solution, the augment will have to work out as x+y=5 3x+y=-1 1 1 5 1 0 1 3 1 -1 to 0 1 4

terenzreignz (terenzreignz):

What were those links for?

OpenStudy (notamathgenius):

The Answers.

OpenStudy (amistre64):

to be more exact ... the "options"

OpenStudy (notamathgenius):

Yea Options

OpenStudy (amistre64):

the solution of Ax = b (A:b) -> ( I : x)

OpenStudy (notamathgenius):

I think I got the "answer" 

OpenStudy (amistre64):

that is not an augment ....

OpenStudy (notamathgenius):

hmm

OpenStudy (notamathgenius):

Let me ask my mom and ill come back if she can't explain it's easier if i can actually see the work being done

OpenStudy (amistre64):

and your links are not effective ....

OpenStudy (notamathgenius):

It's because CA locked them -_-